### v1

# Choose Wisely: Data-Enabled Predictive Control for Nonlinear Systems Using Online Data Selection

Abstract
This paper proposes Select-Data-Enabled Predictive Control (Select-DeePC), a new method for controlling nonlinear systems using output-feedback for which data are available but an explicit model is not. At each timestep, Select-DeePC employs only the most relevant data to implicitly linearize the dynamics in “trajectory space.” Then, taking user-defined output constraints into account, it makes control decisions using a convex optimization. This optimal control is applied in a receding-horizon manner. As the online data-selection is the core of Select-DeePC, we propose and verify both norm-based and manifold-embedding-based selection methods. We evaluate Select-DeePC on three benchmark nonlinear system simulators—rocket-landing, a robotic arm and cart-pole inverted pendulum swing-up—comparing them with standard DeePC and Time-Windowed DeePC, and find that Select-DeePC outperforms both methods. The source code can be found at: https://github.com/naefjo/choose-wisely-paper
Machine Learning, ICML
1 Introduction
Model-based methods have established their value in the Reinforcement Learning (RL) community, even for highly nonlinear systems (Moerland et al., 2022). From a different approach, the Control community has developed interest in “Direct” Data-Driven Predictive Control methods (Dörfler et al., 2023), which typically make Linear, Time-Invariant (LTI) assumptions. This paper proposes Select-Data-enabled Predictive Control (Select-DeePC), a Direct Data-Driven Predictive Control method for nonlinear systems.
Output or safety constraints are challenging for standard RL methods, which learn the constraints by exhaustive trial and error (Sutton & Barto, 1998). Output-constrained systems have traditionally been the domain of Model Predictive Control (MPC) (Morari & H. Lee, 1999), which solves a receding-horizon optimal control problem based on an externally-provided model with an online optimization. In contrast, RL has favored planning methods that are unconstrained and “gradient-free” and, instead of requiring an externally-provided model, use state-action queries of the system. Recently, planning using forward passes through generative or world models have been used to reduce RL sample complexity requirements (Moerland et al., 2022). However, these zero-order-type planning methods usually have no structured way of incorporating output constraints and require the user to offload this complexity into other parts of the method, such as the model itself or the cost.
Figure 1: Closed-loop position trajectories of Select-DeePC (our method), standard DeePC and Time-Windowed DeePC in the MuJoCo Reacher (left) and the Rocket Lander (right) Gym environments, as well as the corresponding open-loop predictions (dashed). In both environments, Select-DeePC converges to the provided setpoint while standard and Time-Windowed DeePC fail to track the setpoints.
Data-Enabled Predictive Control (DeePC) (Coulson et al., 2019) constructs feasible trajectories of an LTI system by linearly combining input-output trajectories in the training data set. DeePC and its variants have demonstrated impressive results on a number of control problems. However, the LTI restriction is severe.
Select-DeePC, proposed in this paper, extends DeePC to the nonlinear setting by iteratively optimizing a convex quadratic program over an implicit linearization in “trajectory space” constructed from data. We define trajectory space as the high-dimensional space in which each dimension corresponds to an input or output over the course of a discrete-time, finite-length trajectory. Linearizing in trajectory space is achieved by selecting only trajectories that are close to the current operating point in trajectory space. Thus, Select-DeePC is a data-driven alternative to nonlinear MPC solved via Sequential Quadratic Programming (SQP), which calculates a trajectory space linearization at each timestep from an externally provided model. SQP-MPC is a standard model-based optimal control tool, commonly used in both academia and industry (Verschueren et al., 2020).
Literature Review DeePC was first introduced by Coulson et al. as an alternative to the sequential, indirect system identification and MPC design pipeline (2019). It uses Willems’ fundamental Lemma (2005) for deterministic LTI systems to generate trajectories of the system as opposed to an identified explicit linear model and is hence considered a direct method. Such predictions fits into the behavioral control framework, see, for instance, (Willems, 1991). For a survey of DeePC we refer the interested reader to (Markovsky et al., 2023).
Several extensions to DeePC have been proposed to handle nonlinear systems in a structured way as opposed to treating the observed nonlinearity as measurement noise on a linear system. Berberich and Allgöwer provide a list of tailored formulations of the Fundamental Lemma to specific classes of nonlinear systems (2024). These formulations enable efficient construction of a DeePC formulation for systems belonging to these specific system classes. However, this already imposes a strong prior on the global nonlinear structure of the dynamical process, which might not be available for an arbitrary system. Alternatively, Berberich et al. propose using a set of the most recent past observations along the current closed-loop trajectory to form the basis of a linear predictor (2022). This predictor is continually updated as new data is obtained in a sliding window fashion, effectively forming a locally linear approximation of the underlying system. As this method only uses recent closed-loop data, it is practical only for systems with benign nonlinearities and with very high signal-to-noise ratios.
Other approaches to nonlinear-DeePC include reformulating the predictor in a kernelized fashion, enabling the use of nonlinear kernels (Lian & Jones, 2021; Huang et al., 2024), using nonlinear basis functions or reformulating the problem in a lifted state-space using Koopman operators (Lazar, 2023). While these methods enjoy greater flexibility due to their use of nonlinear function approximators, the resulting optimization becomes high-dimensional, nonlinear, and generally nonconvex, requiring increased online computational complexity and resulting in suboptimal outcomes.
Figure 2: Venn-diagram demonstrating how the discussed controllers relate and their strengths and weaknesses. Select-DeePC combines the best of all three worlds by bridging the gap of explicit constraint handling without the requirement for Jacobians of the model achieved by direct methods.
The premise of DeePC, solving an optimal control problem based purely on samples of a generative model without requiring gradient information, fits well into the landscape of planning algorithms in model-based RL. Optimization-based planning techniques suffer from accumulating prediction errors (Moerland et al., 2022), exploding gradients due to backpropagation through time or model exploitation in regions of low data density (Kurutach et al., 2018). Currently predominant planning techniques, herein called Sampling-Based MPC, are gradient-free methods that instead rely on sampling of the forward process model, making them computationally cheap to evaluate. Nagabandi et al. propose to generate rollouts of the forward dynamics from random control inputs (2018). Then, the rollouts are scored according to some reward function and the first action of the best trajectory is applied. A more sophisticated version of this same concept is the cross-entropy method for optimization, which uses iterative refinement of the probability distribution over the control inputs (Botev et al., 2013). A related algorithm is Model Predictive Path Integral Control (MPPI) (Williams et al., 2017), which can be interpreted as an importance sampling scheme, weighing the control actions according to their optimality. Sampling-Based MPC in conjunction with RL methods have been successfully demonstrated in various algorithms and applications such as but not limited to (Chua et al., 2018; Hafner et al., 2019; Nagabandi et al., 2019; Kahn et al., 2020). Other recent approaches include using a generative model to generate trajectories and then using an inverse model to predict the required action (Ajay et al., 2023) or simply rely on random shooting and selecting the action sequence with highest reward (Zhou et al., 2024). Figure 2 shows a schematic overview of the control strategies mentioned and their respective strengths and weaknesses.
Remark 1.1.
While we study the performance of Select-DeePC on a fixed data set, which is common in Data-Driven Control literature, and hereby diverge from the common approach in Sampling-Based MPC where a simulator is queried at each decision moment, we note that Select-DeePC can readily be extended to the simulator-based scenario by, e.g., using random roll-outs from the simulator as the data set.
Contributions. We propose Select-DeePC, a data-driven quadratic program predictive controller for nonlinear systems which implicitly linearizes the system dynamics in the extended input-output space by selecting and using only the most relevant data at each decision moment from a fixed data set. It has the following benefits:
  1. 1.
    In contrast with the nonlinear Data-Driven Predictive Control methods in the literature such as basis function methods, Select-DeePC produces trajectories which are consistent with the underlying nonlinear system using a convex quadratic programming solver, yielding a method that can be implemented online in fast dynamic environments.
  2. 2.
    In contrast with model-based RL methods, Select-DeePC allows users to explicitly program-in output constraints and change said constraints, as well as cost/reward function, a posteriori. This allows Select-DeePC to be employed in changing safety conditions without requiring retraining or retuning.
  3. 3.
    Select-DeePC is easy to implement, requiring only a few lines of code for the end-user. Moreover it is conceptually easy to reason about and adapt due to its modular structure. In particular, it enables a rich choice of data-selection methods.
The core idea behind Select-DeePC is the selection of a subset of data from a large data set at each decision moment. We introduce two methods for the data-selection problem, norm-based and manifold-embedding-based selection, and discuss their relevant strengths and weaknesses.
Finally, we demonstrate the benefits of Select-DeePC on three benchmark nonlinear systems including a planar rocket-landing, a two degree of freedom robotic arm, and a cart-pole inverted pendulum, comparing it to standard DeePC and a Time-Windowed DeePC approach (Berberich et al., 2022) and find that Select-DeePC outperforms both in terms of closed-loop cost and prediction accuracy.
2 Select-DeePC
Select-DeePC is a receding-horizon optimal control approach which determines inputs at each timestep using a predictive optimization. While it does not use a model in the classical sense, in the RL parlance, it could be interpreted as a “model-based” method (Sutton & Barto, 1998), while in the Data-Driven Control community it is a “direct” method (Dörfler et al., 2023).
Select-DeePC derives its name from DeePC which solves a receding-horizon predictive control problem subject to input and output constraints. The novelty of DeePC is that it achieves predictive control purely based on implicit predictions of input-output data without ever constructing a explicit predictor (a model) of future measurements given a set of inputs. This is achieved by linearly combining input-output trajectories $\tau$ of length $L$ from a data set ${\mathcal{D}=\{\tau_{0},\dots,\tau_{n_{d}}\}}$. The $L$-length trajectories $\tau_{i}$ are partitioned into a past trajectory $\tau_{\text{p},i}$ and a future trajectory $\tau_{\text{f},i}$ with lengths $T_{\text{p}}$ and $T_{\text{f}}$ respectively with $T_{\text{p}}+T_{\text{f}}=L$. The implicit predictor in DeePC works by pattern matching the past $T_{\text{p}}$ input-output measurements of the current closed-loop trajectory to a linear combination of the $\tau_{\text{p},i}$ in $\mathcal{D}$. Since this can be achieved by (potentially) many different linear combinations, the objective of DeePC is to then find a linear combination of future trajectories $\tau_{f}$ that minimizes a cost function subject to the pattern matching constraint. Several regularizing terms have been introduced to further bias the choice of said linear combinations. For more details, we refer to Appendix A.
Solving DeePC online using a large data set of trajectories is intractable as the solve time of the optimization problem scales superlinearly with the cardinality of the data set. Furthermore, in the presence of nonlinearities, the LTI assumption of DeePC is violated, resulting in suboptimal performance. We now introduce Select-DeePC, an algorithm that allows us to solve a nonlinear data-driven receding-horizon optimal control problem using a convex quadratic programming routine by choosing, at each solver iteration, a subset of trajectories $\tilde{\mathcal{D}}$ with cardinality $\operatorname{card}(\tilde{\mathcal{D}})=N_{\text{cols}}$ from $\mathcal{D}$ where $N_{\text{cols}}\ll\operatorname{card}\left(\mathcal{D}\right)$.
Algorithm 1 Select-DeePC
1:  function Select-DeePC($u_{\text{p}}$, $y_{\text{p}}$
2:     while convergence criterion not met do
[…]
4:        $\tilde{D}\leftarrow\texttt{select}(\mathcal{D},\,\tilde{\tau},\,N_{\text{cols}})$
5:        $u\leftarrow\texttt{DeePC.computeAction}(u_{\text{p}},\ y_{\text{p}},\ \tilde{D})$
6:     end while
7:     Return $u_{0}$
8:  end function
Select-DeePC is summarized in Algorithm 1. In between two real-time sampling instances, an iterative procedure is performed where, at every iteration, we select from a data set $\mathcal{D}$ a subset of trajectories $\tilde{\mathcal{D}}$ that are “relevant” to the current linearization point given by the open loop solution $\tilde{\tau}$. While we loosely use the term relevance here, this mechanism will be discussed in more detail in Section 3. Select-DeePC then solves a Data-Driven Predictive Control problem using the subset of data. This is iterated until some convergence criterion is met, e.g. the open-loop prediction $\tilde{\tau}$ converges.
We now discuss a number of attributes of this proposed method.
Cost-agnostic. Since the training process/data set generation is structurally independent of the control formulation, a posteriori cost or constraint modification is possible, allowing the method to be applied zero-shot to new problems without the need for retraining.
Hyperarameters. The performance of Select-DeePC depends on the tuning of several hyperparameters. First, the DeePC specific hyperparameters, such as regularization terms and horizon lengths $T_{\text{p}},\ T_{\text{f}}$. These can either be hand-tuned according to established heuristics (Elokda et al., 2021) or optimized through gradient-descent methods (Cummins et al., 2024) or black-box parameter tuning (Berkenkamp et al., 2016). For DeePC tuning guidelines, we refer to (Markovsky et al., 2023). Select-DeePC adds onto these the design parameters $N_{\text{cols}}$ as well as the method of data selection. In the following section, we will investigate two different data-selection methods which subsample a specified number of trajectories from a large offline library of trajectories observed during “training.”
SQP-MPC. The iterative refinement of the solution until convergence is inspired by the model-based SQP-MPC algorithm for nonlinear systems. We direct the interested reader to Appendix B for a detailed discussion of SQP-MPC and the relationship to Select-DeePC. We note that Select-DeePC’s implicit linearization is performed in trajectory space. In the model-based case, linearizing in trajectory space corresponds to a linear time-varying parametrization of the nonlinear dynamics for the MPC. This linear, time-varying parametrization contrasts with an LTI formulation in which the Jacobians are only evaluated at $y(k)$ and kept fixed along the prediction horizon.
Computational complexity. Solving the convex quadratic optimization problem at the basis of Select-DeePC involves repeatedly solving a symmetric system of linear equations (Stellato et al., 2020). Hence, we expect the algorithm to scale with $\mathcal{O}(n^{3})$ where $n$ indicates the number of optimization variables which is directly proportional to the number of trajectories $N_{\text{cols}}$ selected from the full data set. Hence, from a computational complexity standpoint, we wish to perform the optimization with as few trajectories as possible. Additionally, the dominant computation in the data-selection is the sorting of the data according to the relevance to $\tilde{\tau}$ which scales with $\mathcal{O}(n_{d}\log n_{d})$, which we postulate is less significant than the computational complexity of the optimization problem. While Select-DeePC solves a convex quadratic program and is smaller than full-data DeePC or nonlinear DeePC, it still requires more compute than Sampling-Based MPC methods. There exist DeePC variants which reduce the number of optimization variables in the predictive control problem but the comparison of these methods is outside the scope of this study (Favoreel & De Moor, 1999; Breschi et al., 2023). For an overview of existing methods we refer to (Verheijen et al., 2023).
3 Data Selection
At the heart of Select-DeePC lies the data selection method, which selects data points (representing trajectories $\tau_{i}$) from a large (offline) data set. Since the subproblem solved by Select-DeePC relies on linearly combining trajectories, the subset of data should exhibit predominantly linear dynamics. The suggested approach in the literature is to use a sliding window over past input-output measurements. We maintain the view that $\tilde{\mathcal{D}}$ should be chosen according to some spatial distance metric (in trajectory space) as opposed to temporal proximity (such as a sliding window) to the current linearization point $\tilde{\tau}$.
In the following, we will present two selection methods, norm-based and manifold-embedding-based selection, while placing emphasis on the modular nature of Select-DeePC. While norm-based selection is efficient to implement and allows the user to easily update the data set online, manifold-embedding-based selection enables the distance computation in a lower dimensional embedding space which lessens the impact of the curse of dimensionality but comes with increased offline computation cost.
3.1 Norm-based selection
Norm-based data selection selects from a data set $\mathcal{D}$ the $N_{\text{cols}}$ closest (in space) trajectories by computing the norm of the differences between the open-loop solution $\tilde{\tau}$ and the trajectories in the data set $\mathcal{D}$. An implementation is given in Algorithm 2
Algorithm 2 Norm-based Data Selection
1:  function NormDataSelection($\mathcal{D}$, $\tilde{\tau}$, $N_{\text{cols}}$
2:     $\tilde{\mathcal{D}}$ = sort($\left(\{\|\tau_{i}-\tilde{\tau}\|\,|\,i=1,\dots,n_{d}\}\right)$
3:     Pick first $[1,N_{\text{cols}}]$ from $\tilde{\mathcal{D}}$
4:  end function
While this approach is intuitive, it is common knowledge that norm-based distances fall flat in high dimensional spaces due to the curse of dimensionality (Aggarwal et al., 2001). Since we are computing norm-based distances of trajectories, i.e. vectors with dimension $(T_{\text{p}}+T_{\text{f}})(m+p)$, this method is not an exception to that.
3.2 Low Dimensional Representation using Manifold Learning
The field of Manifold Learning (Nonlinear Dimensionality Reduction) concerns itself with the compression of high-dimensional data onto manifolds with lower intrinsic dimension than the original data by using (nonlinear) projection techniques (Lee & Verleysen, 2007; Jia et al., 2022). Manifold Learning techniques rely on the manifold hypothesis (Fefferman et al., 2016), which postulates that the intrinsic dimensionality of a data set $\mathcal{D}$ is much lower than the dimensionality of the individual data points. For instance, if the system under consideration is a controllable and deterministic LTI system, then Willems’ Fundamental Lemma states that the trajectories span a manifold (in particular a subspace) that can at most have $n+T_{\text{f}}\cdot m$ degrees of freedom (2005), which is generally less than the trajectory dimensionality $(T_{\text{p}}+T_{\text{f}})(m+p)$. Here, $n$ denotes the dimension of the processes latent state, $p$ is the dimension of the measurement vector and $m$ indicates the dimension of the input signal. Motivated by this observation, we wish to find a lower dimensional representation of the data where the similarities can be computed without incurring the curse of dimensionality.
Isomap, introduced by Tenenbaum et al. (2000), is an unsupervised manifold learning method which aims to find an embedding of the data in a Euclidean space which preserves geodesic distances of a neighborhood graph of the data constructed in the data space. This comes with increased computational cost compared to norm based data selection but can offer inter data point distances which more accurately reflect the nonlinear nature of the data manifold in trajectory space. Also notice that the bulk of this added cost is a one time upfront cost (computing the embedding), which can be performed offline. For algorithmic details, see Appendix C.
By considering the reconstruction error of the data, an estimate of the dimensionality of the data can be found. This dimensionality can be interpreted as the intrinsic dimensionality of the underlying system. See Appendix C for plots demonstrating the diminishing returns of higher embedding dimensions. The relative contrast of a data set $\Delta=(d_{\text{max}}-d_{\text{min}})/d_{\text{min}}$, with $d_{\text{max/min}}$ indicating the maximum and minimum norm values of the data set respectively, is a helpful surrogate to determine how distinct each data point appears under a given distance metric. Figure 3 shows the relative contrast of a data set collected from the rocket simulator in Section 4 as a function of trajectory dimensionality when embedded using Isomap. It compares relative contrast given a fixed embedding dimensionality as well as a varying one, which increases according to $n+T_{\text{f}}\cdot m$, to the relative contrast of the $L_{1}$ norm. While a varying embedding dimensionality offers greater relative contrast for small trajectory lengths, we encounter diminishing returns for large trajectory lengths.
Figure 3: Relative contrast with Isomap embedding compared to $L_{1}$ selection. Both Isomap methods have better relative contrast compared to $L_{1}$ selection. Furthermore, we observe that increasing the embedding dimension further leads to diminishing returns in relative contrast compared to a fixed embedding dimension.
Although this does not fully resolve the problem caused by the curse of dimensionality as the embedding dimension, i.e. the dimension where we need to compute distances, will still scale with at least with $n+T_{\text{f}}\cdot m$, it does offer a structured way of compressing the data to a reasonable dimension. Experiments showed that embedding dimensions lower than the minimum dimension required to represent a linear system already lead to diminishing returns in reconstruction error.
Algorithm 3 Manifold-Embedding-based Data Selection
1:  function ManifoldDataSelection($\mathcal{D}$, $\tilde{\tau}$, $N_{\text{cols}}$
2:     $\mathcal{D}_{\text{e}}\leftarrow\texttt{embed}\left(\mathcal{D}\right)$
3:     $\tilde{\tau}_{\text{e}}\leftarrow\texttt{embed}\left(\tilde{\tau}\right)$
4:     $\tilde{\mathcal{D}_{\text{e}}}$ = sort($\left(\{\|\tau_{\text{e},i}-\tilde{\tau}_{\text{e}}\|\,|\,i=1,\dots,n_{d}\}\right)$
5:     Pick first $[1,N_{\text{cols}}]$ from $\tilde{\mathcal{D}}_{\text{e}}$
6:  end function
The pseudocode for manifold-embedding-based data selection using the Isomap algorithm is outlined in Algorithm 3, the details of embed using Isomap can be found in Appendix C. Notice that the embedding of the data set in line 2 can be cached/computed offline. Online, the algorithm then proceeds to embed the query trajectory $\tilde{\tau}_{i}$ and computes its similarities in the lower dimensional embedding space. Then, similarly to Algorithm 2, the relevant trajectories are returned in the trajectory space, after which the implicit predictor for DeePC is constructed using the retrieved non-embedded data points.
4 Results and Discussion
Simulation experiments were performed to validate the efficacy of Select-DeePC and to benchmark it against classical DeePC, which utilizes the entire data set, as well as the Time-Windowed DeePC variant proposed by Berberich et al. (2022), which constructs the implicit predictor by continuously updating the data set with new input-output data and discarding the most out-of-date measurement.
4.1 Landing a Reusable Rocket
We use a planar vertical takeoff, vertical landing rocket to demonstrate the performance of Select-DeePC on an open-loop unstable nonlinear system. We show that Select-DeePC successfully copes with the nonlinear dynamics and manages to stabilize the system at a given setpoint. A custom environment [^1] https://gitlab.ethz.ch/bsaverio/coco-project for the gymnasium suite (Kwiatkowski et al., 2024) was used as a simulation platform. The rocket measurement $y\in\mathbb{R}^{6}$ includes the 2 dimensional position, the corresponding velocities, the heading, and the angular rate of the rocket. The rocket control input $u\in\mathbb{R}^{3}$ is the main engine thrust, the angle of the main engine gimbal, and the thrust magnitude of the side thrusters. For a detailed description of the environment, we refer to (Cummins et al., 2024).
The right column in Figure 1 shows the simulator setup and specifically, in the bottom figure, three closed-loop rocket trajectories. One for Select-DeePC, one for standard DeePC and one for Time-Windowed DeePC respectively. While both the standard DeePC and the Time-Windowed DeePC controllers diverge, Select-DeePC manages to regulate the system to the desired setpoint and stabilize it there.
Ablation Study. We compare the performance of Select-DeePC using the two presented selection methods on two different data sets. Both data sets were collected using input sequences of the form
$u_{t}=a\cdot u_{t-1}+b\cdot n,\,n\sim\mathcal{U}(-1,1).$ (1)
The data set IID Data uses $(a,b)=(0,1)$ and Random Walk Data uses $(a,b)=(1,0.1)$. To ensure that a large region of the trajectory space was covered by the collected data, 100 simulations were run with a maximum of 100 timesteps each, resulting in data sets with $5000$$ data points. As shown later, the level of nonlinearity in the data has a significant effect on the performance of standard DeePC. Due to the high inertia of the rocket, the IID Data predominantly exhibits behaviors that can be represented reasonably well by a linear model, and hence, standard DeePC can be applied. However, this comes at a performance penality as nonlinear effects are not accounted for. We observe that Select-DeePC still outperforms standard DeePC on IID Data and that data selection can still lead to performance increases if the data set does not fully capture the nonlinear behavior of the underlying system. The random walk structure of the input sequence in Random Walk Data, on the other hand, ensures that the nonlinearities of the system are sufficiently excited. This makes standard DeePC fail but provides Select-DeePC with more global information about the nonlinear system.
The comparisons of Select-DeePC on the two data sets are two-fold. First, the resulting predictive accuracy of the selection methods are compared. Furthermore, a comparison of the realized closed-loop cost is given as a function of selection method and data set. As a benchmark, random sketching of the full data set is used at each decision moment.
Figure 4: Residual between the ground-truth and the least-squares prediction as a function of trajectories used in the predictor. The selection methods compared here are using random sketching of the data set, $L_{1}$-norm based and Isomap-embedding based selection. Notice that data selection leads to lower cumulative residual than using the full data set (right most data point).
The predictive accuracy of the methods was compared by computing the least squares solution of the predictor used in the DeePC subproblem as a function of the number $N_{\text{cols}}$ of trajectories in the selected set of data. For each trajectory in the holdout data set, constructed from trajectory-segments of a closed-loop trajectory, the $N_{\text{cols}}$ closest trajectories (in spatial $L_{1}$-norm or using Isomap-embedding) under the given selection method were selected and the prediction of future measurements $\hat{y}_{\text{f}}$, given the corresponding future input sequence $u_{\text{f}}$ was computed using Equation 11. The plots in Figure 4 show the cumulative residuals between the predicted future trajectory and the ground truth realization $\|\hat{y}_{\text{f}}-y_{\text{f}}\|_{2}$ over the entire validation set.
In the extreme case where $N_{\text{cols}}=\operatorname{card}(\mathcal{D})$, each of the methods coincides with the least squares solution of the predictor in standard DeePC. We observe that both our methods of data selection (based on spatial $L_{1}$-norm or Isomap-embedding) lead to lower cumulative residual on the holdout validation set when compared to using the full data set or random sketching. In particular, there is a sweet spot in the number $N_{\text{cols}}$ of trajectories to use in the predictor.
Figure 5: Closed-loop cost as a function of data subset cardinality on two different data sets. Instances where the selection methods realized infinite cost (due to instability) have been omitted and the remaining finite cost data points are shown. Selecting data points outperforms using the full data on all data sets.
Figure 5 shows the closed-loop performance of Select-DeePC on the setpoint tracking task depicted in the right column of Figure 1 on both IID Data and Random Walk Data. We again observe on both data sets that selecting data outperforms using the full data set in terms of closed-loop performance, as a lower cost could be achieved. Using IID Data, the reduction in closed-loop cost is 1.9 while using Random Walk Data, DeePC using the full data set failed to stabilize the rocket.
Coincidentally, we observe that random sketching of the data set only results in marginally worse performance compared to Isomap or $L_{1}$ selection. This could be explained by the fact that the system is still relatively close to linear as the attitude of the rocket does not exceed 60°60°60  ° and the high inertia of the system can help in compensating bad actions taken in one timestep—as the randomly selected data are resampled every timestep, it is unlikely that the sampled data poorly represents the system dynamics in multiple consecutive timesteps.
Computational Cost.
Figure 6: Solve times of Select-DeePC as a function of number of Hankel columns decomposed into data selection time and QP solve time. While the QP solve time increases as the number of Hankel columns increases, the data selection time stays constant. It is only affected by the size of the entire data set which stays constant.
Figure 6 shows the computation times of Select-DeePC as a function of number of subselected trajectories. As postulated in Section 2, the dominant factor in the solve times is solving the quadratic program that scales at least quadratically with $N_{\text{cols}}$, while data selection is a constant time operation with respect to $N_{\text{cols}}$.
4.2 Planar Robotic Manipulator
We use the Reacher (Kwiatkowski et al., 2024) environment, which simulates a planar robotic manipulator, to demonstrate the ability of Select-DeePC to control a highly nonlinear system and successfully track reference setpoints. Furthermore, we show that Select-DeePC is effective in coping with and respecting a posteriori specified output constraints. The kinematic chain of the manipulator consists of two rigid bodies that are connected by revolute joints which are each actuated by a torque, giving it an action space of $u\in\mathbb{R}^{2}$. The measurement vector consists of the sine and cosine of both joint angles respectively, and the end-effector position in $\mathbb{R}^{2}$. Furthermore, we are also able to measure the angular velocity of each link, resulting in an observation space $y\in\mathbb{R}^{8}$. This system is more challenging to control using DeePC (or any linear control method) since the end-effector position is dictated by the composition of two rotations. In this environment, it is not unreasonable to assume that the angles must exceed $90^{\circ}$ in order to reach a target. Furthermore, for a given end-effector position, the corresponding joint configuration is not necessarily unique.
Ablation Study. We again investigate the effect of number of selected trajectories $N_{\text{cols}}$, as well as selection strategy on the resulting closed loop cost. For the Reacher, data were collected using IID inputs. In total, 200 simulations were run with 200 steps each. Here, only IID inputs were collected instead of random walk because the system has lower inertia compared to the rocket and the nonlinearities could be excited more easily.
[…]
Figure 7: Comparing the closed-loop cost incurred as a function of subset cardinality. Clearly there is a sweet spot in how many trajectories should be selected for the predictions. Furthermore, due to the highly nonlinear nature of the problem, distance based data selection clearly outperforms random sketching of the full data set.
Figure 7 again shows the incurred closed-loop cost as a function of number of trajectories used in the implicit predictor. Similarly to the rocket, we observe a decrease in predictive accuracy and, as a consequence, an increase in cost as the number of data points used in the predictor increases. While random sketching of the data set showed decent results in the rocket simulation, it fails completely in the reacher simulation as no accurate predictions could be generated, resulting in the controller not making any progress towards the setpoint and hence accumulating high closed-loop cost.
Figure 8: Closed-loop trajectories for a constrained (dark blue) and unconstrained (blue) DeePC subproblem.
Zero-Shot Generalizability of Select-DeePC. Figure 8 shows two closed-loop trajectories of Select-DeePC where the controller was tasked with tracking a set of consecutive end-effector references. We observe that Select-DeePC is able to generalize to different tasks across the entire trajectory manifold of the Reacher by reusing the same data set and without structural changes to the controller. While one simulation was left unconstrained, a constraint of $y<0.1$ was added to the DeePC subproblem of the other one, indicated by the red region. The structure of Select-DeePC effectively decouples the dynamics learning/training process from the controller design. Indeed, both the constrained and unconstrained tasks use the same data set which is unaware of the structure of the predictive control problem. This allows us to easily adapt the cost function or constraint terms for the task at hand in zero-shot fashion without requiring task specific data or additional online learning.
4.3 Cart-Pole inverted pendulum swing-up
Figure 9: Closed-loop trajectory for an inverted pendulum swing-up. Select-DeePC (blue) is successful in transitioning from the downward facing stable equilibrium to the upward facing unstable equilibrium. Standard DeePC (orange) on the other hand fails to transition between the equilibria. The dashed blue line indicates the open-loop prediction of Select-DeePC at the four snapshot time instances, t=15, 22, 32 and 120.
Finally, we show that Select-DeePC is capable of performing a cart-pole inverted pendulum swing-up, which is a standard benchmark for nonlinear MPC methods. This test demonstrates that Select-DeePC successfully plans trajectories for the nonlinear dynamics using its implicit linearization in trajectory-space. The input $u\in\mathbb{R}$ of the inverted pendulum is a force acting on the cart along the $x$-axis. Furthermore, the measurement $y\in\mathbb{R}^{5}$ consists of the cart position, the sine and cosine of the angle between the pendulum and the vertical axis, as well as linear and angular velocities of the cart and pendulum respectively.
Due to the unstable nature of the desired operating point and the system starting in a stable equilibrium position, data was gathered in closed-loop. Specifically, the data set consists of 200 demonstrations of successful but suboptimal swing-ups, collected using an energy based feedback controller (Chatterjee et al., 2002).
Figure 9 shows the closed-loop trajectories of both Select-DeePC and standard DeePC when tasked with tracking a setpoint at the upward facing, unstable equilibrium while starting from the downward facing stable equilibrium. While standard DeePC is unable to perform the swing-up and oscillates around the stable equilibrium, Select-DeePC is successfully able to learn from the demonstrations in the data set and performs a swing-up to the desired vertical position and subsequently stabilizes the pendulum there. It is worthwhile to point out that the controller does not simply select the top trajectory in the selected data set. Instead, Select-DeePC linearly combines several trajectory segments at each iteration of the solver to produce the swing-up (c.f. Figure 10). This is also evident from the fact that the horizon in the predictive controller is shorter than the episode length, thus the solver never has access to a full swing-up trajectory.
Figure 10: Number of trajectories used to construct the final prediction at each sampling instant of the inverted pendulum swing-up.
5 Conclusion
This paper introduces Select-DeePC, a novel approach to constraint-aware Data-Driven Predictive Control for nonlinear systems. At each time instance, the method pre-processes the data set to determine the most relevant data, then passes the most relevant data to a convex optimization which determines the optimal control for the future horizon. The optimal control for the given timestep is implemented, and the algorithm progresses in a receding horizon fashion.
In contrast to other DDPC methods, our method can be implemented easily, scales favorably with the number of collected data-points and increases performance outside of the linear domain.
We validated this in a variety of forms and on a set of three simulation environments. In the rocket simulator environment, we showed that for one of the data-sets considered, Select-DeePC cut closed-loop cost by a factor of 2. For the other data-set, our method succeeded, where previous methods crashed. In the second set of simulations in the Reacher environment, we enabled zero-shot constrained setpoint tracking of a reference signal in a robotic reacher simulation. Again, this was not achievable using previous DeePC formulations. Finally, in the last set of simulations of a cart-pole inverted pendulum, we demonstrated successful transitioning between two different equilibria of the system by performing a pendulum swing-up.
A direction of future work is to investigate different selection methods, such as structure-exploiting selection methods or selection methods based on an information criterion. Another direction is to replace Select-DeePC’s large training data set with a data-driven forward simulator of the dynamics (i.e., a “world model”), resulting in a constraint-aware alternative to Sampling-Based MPC.
Acknowledgments
The authors would like to thank Xian Li for providing the simulation environment and data collection procedure for the inverted pendulum.
Appendix A DeePC
Data-enabled Predictive Control (DeePC) solves a receding horizon optimal control problem based purely on data for a linear time-invariant system (LTI) of the form
$y_{\text{f}}(k)=F_{\text{p}}\begin{bmatrix}u_{\text{p}}(k-1)\\ y_{\text{p}}(k-1)\end{bmatrix}+F_{\text{f}}u_{\text{f}}(k),$ (2)
where $u$ and $y$ are partitioned into “past” and “future” quantities $u_{\text{p}},u_{\text{f}}$ and $y_{\text{p}},y_{\text{f}}$ respectively. This means that the sequence of predicted future measurements, the past measurements used by the predictor, the sequence of past inputs applied to the system and the future control input sequence respectively given by
$y_{\text{f}}(k)$ $\coloneqq\begin{bmatrix}y(k)^{\top}&\dots&y(k+T_{\text{f}}-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{T_{\text{f}}\cdot p},$
$u_{\text{f}}(k)$ $\coloneqq\begin{bmatrix}u(k)^{\top}&\dots&u(k+T_{\text{f}}-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{T_{\text{f}}\cdot m},$
$y_{\text{p}}(k-1)$ $\coloneqq\begin{bmatrix}y(k-T_{\text{p}})^{\top}&\dots&y(k-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{T_{\text{p}}\cdot p}\text{, and}$
$u_{\text{p}}(k-1)$ $\coloneqq\begin{bmatrix}u(k-T_{\text{p}})^{\top}&\dots&u(k-1)^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{T_{\text{p}}\cdot m}.$
Let $\mathcal{U}$ and $\mathcal{Y}$ denote constraint sets on the input and output $u_{\text{f}}$ and $y_{\text{f}}$ respectively. Given a set of stage cost functions $c_{i}$ and past trajectory of the system, we are interested in resolving the following problem in a receding horizon fashion.
$\min_{u_{\text{f}},y_{\text{f}}}\quad$ $\sum_{i=0}^{T_{\text{f}}-1}c_{i}(u_{\text{f},i},y_{\text{f},i})$ (3a)
s.t. $y_{\text{f}}=F_{\text{p}}\begin{bmatrix}u_{\text{p}}\\ y_{\text{p}}\end{bmatrix}+F_{\text{f}}u_{\text{f}},$ (3b)
$(u_{\text{f}},y_{\text{f}})\in\mathcal{{U}\times\mathcal{Y}}.$ (3c)
Without access to the model (2), we will need to reconstruct it from offline data. Let $v_{T}=\left\{v(i)\right\}_{i=0}^{T-1}$ denote measurements of a signal $v$ with length $T\in\mathbb{Z}_{\geq 1}$. We define the Hankel matrix of depth $L$ of $v_{T}$ as
$H_{L}(v_{T})=\begin{bmatrix}v(0)&v(1)&\dots&v(T-L-1)\\ v(1)&v(2)&\dots&v(T-L)\\ \vdots&\vdots&\ddots&\vdots\\ v(L-1)&v(L)&\dots&v(T-1)\end{bmatrix}$ (4)
Note that each column of the Hankel matrix contains a trajectory of length $L$.
Definition A.1 (Persistency of Excitation).
For a given sequence $v_{T}$, we call the sequence persistently exciting of order $L$ if $H_{L}(v_{T})$ has full row rank.
Using persistently exciting inputs, Willems’ Fundamental Lemma allows us to parametrize all finite-length trajectories of a controllable linear system.
Lemma A.2 (The Fundamental Lemma (Willems et al., 2005)).
Consider a controllable linear time-invariant system of order $n$. Given an input sequence $u_{T}$ which is persistently exciting of order $L\cdot m+n$ and the corresponding output sequence $y_{T}$, then there exists $g$ such that any admissible trajectory (u, y) of length $L$ can be expressed as a linear combination
$\begin{bmatrix}H_{L}(u_{T})\\ H_{L}(y_{T})\end{bmatrix}g=\begin{bmatrix}u\\ y\end{bmatrix}.$ (5)
DeePC is based on the use of this implicit predictor (5) as a proxy for the explicit model representation in (2). In an online manner, DeePC matches the “past” data with the most recently seen inputs and outputs, and uses the “future” data to match a prediction over which we optimize. This results in a receding-horizon predictive controller based purely on data:
$\min_{u_{\text{f}},y_{\text{f}},g}\quad$ $\sum_{i=0}^{T_{\text{f}}-1}c_{i}(u_{\text{f},i},y_{\text{f},i})+r(g)$ (6a)
s.t. $\begin{bmatrix}H_{T_{\text{p}}+T_{\text{f}}}(u_{T})\\ H_{T_{\text{p}}+T_{\text{f}}}(y_{T})\end{bmatrix}g=\begin{bmatrix}u_{\text{p}}\\ u_{\text{f}}\\ y_{\text{p}}\\ y_{\text{f}}\end{bmatrix},$ (6b)
$(u_{\text{f}},y_{\text{f}})\in\mathcal{{U}\times\mathcal{Y}}.$ (6c)
Here, $r(g)$ denotes suitable regularizers on $g$ (Dörfler et al., 2023) which improve robustness of the controller in the presence of noise. The implicit predictor in (5) can readily be extended to affine systems as follows (Martinelli et al., 2022; Berberich et al., 2022)
$\begin{bmatrix}H_{t}(u_{T})\\ H_{t}(y_{T})\\ \mathbbm{1}^{\top}\end{bmatrix}g=\begin{bmatrix}u\\ y\\ 1\end{bmatrix}.$ (7)
A.1 From Implicit to Explicit Predictor
Partitioning the predictor (7) into past and future states results in
$\begin{bmatrix}U_{\text{p}}\\ U_{\text{f}}\\ Y_{\text{p}}\\ Y_{\text{f}}\\ \mathbbm{1}^{\top}\end{bmatrix}g=\begin{bmatrix}u_{\text{p}}\\ u_{\text{f}}\\ y_{\text{p}}\\ y_{\text{f}}\\ 1\end{bmatrix},$ (8)
which is an implicit predictor for the future measurement trajectory $y_{\text{f}}$ given past measurements $y_{\text{p}}$, the corresponding past input sequence $u_{\text{p}}$ and a future input sequence $u_{\text{f}}$. Using the partitions
$H_{z}\coloneqq\begin{bmatrix}U_{\text{p}}\\ U_{\text{f}}\\ Y_{\text{p}}\\ \mathbbm{1}^{\top}\end{bmatrix},\ z\coloneqq\begin{bmatrix}u_{\text{p}}\\ u_{\text{f}}\\ y_{\text{p}}\\ 1\end{bmatrix},$ (9)
we can eliminate $g$ from the implicit predictor in Equation 8 and write the equivalent explicit predictor
$y_{\text{f}}=Y_{\text{f}}H_{z}^{\dagger}z+Y_{\text{f}}N_{H_{z}}g_{0},$ (10)
where $A^{\dagger}$ denotes the pseudo-inverse of $A$, $N_{H_{z}}$ is the null-space projection matrix of $H_{z}$ and $g_{0}$ is an arbitrary perturbance to the least squares solution $g=H_{z}^{\dagger}z$. Assuming $H_{z}$ has full row rank, then the least squares solution (i.e. $N_{H_{z}}g_{0}=0$) is
$y_{\text{f}}=Y_{\text{f}}H_{z}^{\top}\left(H_{z}H_{z}^{\top}\right)^{-1}z.$ (11)
Appendix B SQP-MPC
Nonlinear MPC tries to optimize a cost function subject to the state evolution of a nonlinear system of the form
$y_{\text{f}}(k)=f\left(u_{\text{p}}(k-1),u_{\text{f}}(k),y_{\text{p}}(k-1)\right).$ (12)
If we had access to $f$, then we could use the receding-horizon optimal control problem
$\min_{u_{\text{f}},y_{\text{f}}}\quad$ $\sum_{i=0}^{T_{\text{f}}-1}c_{i}(u_{\text{f},i},y_{\text{f},i})$ (13a)
s.t. $y_{\text{f}}=f\left(u_{\text{p}},u_{\text{f}},y_{\text{p}}\right),$ (13b)
$(u_{\text{f}},y_{\text{f}})\in\mathcal{{U}\times\mathcal{Y}}$ (13c)
to obtain an optimal control input sequence $u^{\star}_{\text{f}}$. One method of solving a nonlinear optimal control problem of this form is using SQP, which repeatedly solves a linearization of the nonlinear problem (13a), resulting in subproblem (14). The solution of said subproblem is then used to update the estimate of the optimal open-loop solution $(\tilde{u}_{\text{f}},\tilde{y}_{\text{f}})$.
$\min_{\Delta u_{\text{f}},\Delta y_{\text{f}}}\quad$ $\sum_{i=0}^{T_{\text{f}}-1}\begin{bmatrix}\Delta y_{\text{f},i}^{\top}&\Delta u_{\text{f},i}^{\top}&1\end{bmatrix}\tilde{H}_{i}\begin{bmatrix}\Delta y_{\text{f},i}\\ \Delta u_{\text{f},i}\\ 1\end{bmatrix}$ (14a)
s.t. $\Delta y_{\text{f}}=F_{\text{p}}\begin{bmatrix}\Delta u_{\text{p}}\\ \Delta y_{\text{p}}\end{bmatrix}+F_{\text{f}}\Delta u_{\text{f}},$ (14b)
$\Delta z_{\text{p}}=\begin{bmatrix}u_{\text{p}}(k)-\tilde{u}_{\text{p}}\\ y_{\text{p}}(k)-\tilde{y}_{\text{p}}\end{bmatrix},$ (14c)
$(\Delta u_{\text{f},i},\Delta y_{\text{f},i})\in\mathcal{U}\times\mathcal{Y}\quad\forall i\in\left\{0,\dots,T_{\text{f}}-1\right\},$ (14d)
where the symbols $F_{\text{p}}\coloneqq\frac{\partial}{\partial\begin{bmatrix}u_{\text{p}}^{\top}&y_{\text{p}}^{\top}\end{bmatrix}^{\top}}f(u_{\text{p}},y_{\text{p}},\tilde{u}_{\text{f}})$ and ${F_{\text{f}}\coloneqq\frac{\partial}{\partial\tilde{u}_{\text{f}}}f(u_{\text{p}},y_{\text{p}},\tilde{u}_{\text{f}})}$ indicate the respective Jacobians of system (12) evaluated at the current solution estimate $(u_{\text{p}},y_{\text{p}},\tilde{u}_{\text{f}},\tilde{y}_{\text{f}})$.
The full SQP algorithm can be summarized as follows
Algorithm 4 SQP-MPC (IO Representation)
1:  function SQP-MPC($z_{\text{p}}$
2:     while not converged do
3:        linearize dynamics, constraints and cost at $(z_{\text{p}},\tilde{u}_{\text{f}},\tilde{y}_{\text{f}})$
4:        Solve Equation 14
5:        $(\tilde{u}_{\text{f}},\tilde{y}_{\text{f}})\leftarrow(\tilde{u}_{\text{f}},\tilde{y}_{\text{f}})+(\Delta u_{\text{f}},\Delta y_{\text{f}})$
6:     end while
7:     Return $\tilde{u}_{\text{f},0}$
8:  end function
B.1 Equivalence between Select-DeePC and SQP-MPC
Select-DeePC has a very natural interpretation as solving a nonlinear MPC problem using SQP, where the analytical Jacobian computation is replaced by a data-driven approach of estimating the Jacobian of the input-output behavior from data. Specifically, one can select data points “close” to the current linearitation point (given by the open loop solution), resulting in an equivalent, data-driven approximation of $F_{\text{p}}$ and $F_{\text{f}}$ in Equation 14 that approaches the analytical Jacobians assuming sufficient amounts of data in the neighbourhood around the linearization point (i.e. $\operatorname{card}(\mathcal{D})$ tending to infinity).
Appendix C Isomap
Isomap, short for Isometric Mapping, is a nonlinear dimensionality reduction technique introduced by Tenenbaum (2000). It can be seen as an extension of classical Multidimensional Scaling (MDS) where, instead of using Euclidean interpoint distances, geodesic distances between the data points are preserved. These geodesic distances are estimated using shortest distances on a connected graph which connects neighboring data points. These shortest paths are then arranged in a distance matrix $D$. Isomap then finds an embedding by applying classical MDS. Computing the distance of a new query data point to the data set in the embedding dimension is done by first linking the new data point into the neighborhood graph and then projecting the data point into the embedding space.
Isomap is able to capture the overall nonlinear nature of the data manifold and is guaranteed to produce a global optimizer of its reconstruction loss. Furthermore, due to preserving geodesic distances, points close in the input space are also close in the embedding space, resulting in interpretable results. This is especially relevant for Select-DeePC which is build upon the idea of selecting data points according to their associated distance to a query point.
C.1 Isomap Parameter Selection
Isomap comes with a set of hyper parameters that need to be selected for the data set at hand. Specifically, the number of closest data points that should be considered as neighbors during the neighborhood graph construction and the dimensionality of the embedding space. Figure 11 shows the reconstruction error of the data sets as a function of number of graph neighbors used during adjacency graph construction and number of dimensions of the embedding space. The graphs show a minimum around 10 graph neighbors while we can observe diminishing returns from increasing the embedding dimensionality from 64 to 128 which is consistent with the formula $T_{\text{f}}\cdot m+n$ which suggests that the locally linear system has a dimensionality of 96.
Furthermore, note the sharp increase in reconstruction error as the number of graph neighbors increases after the local minimum. This demonstrates the effect of bad hyperparameter selection which can lead to short-circuiting or over-connectedness in the graph, leading to worse embedding performance as the local manifold structure is lost (Tenenbaum et al., 2000).
Figure 11: Isomap reconstruction error as a function a embedding dimensionality and number of neighbors used in the adjacency graph.
Appendix D Experiment Hyperparameters
All experiments were performed using a quadratic tracking cost of the form
$c_{i}(u_{\text{f},i},y_{\text{f},i})=\|y_{\text{f},i}-y_{r}\|_{Q}+\|u_{\text{f},i}\|_{R},$ (15)
where $Q$ and $R$ are positive (semi-)definite cost matrices and $y_{r}$ is a reference setpoint. Furthermore, for the regularization term $r(g)$ we use 1-norm and projection regularization (Markovsky et al., 2023) with weights $\lambda_{1}$ and $\lambda_{\Pi}$ respectively. The following tables shows the hyperparameters for the three sets of simulations.
Parameter Value
$Q$ $\operatorname{diag}\left(\begin{bmatrix}40&20&20&1&3000&30\end{bmatrix}\right)$
$R$ $\operatorname{diag}\left(\begin{bmatrix}10&10&10\end{bmatrix}\right)$
$\lambda_{1}$ 0
$\lambda_{\pi}$ 5000
Table 1: Hyperparameters for the rocket simulation environment.
Parameter Value
[…]
$R$ $\operatorname{diag}\left(\begin{bmatrix}10&10\end{bmatrix}\right)$
$\lambda_{1}$ 10
$\lambda_{\pi}$ 10 0001000010\,00010 000
Table 2: Hyperparameters for the reacher simulation environment.
Parameter Value
[…]
$R$ $1$
$\lambda_{1}$ 50 0005000050\,00050 000
$\lambda_{\pi}$ 0
Table 3: Hyperparameters for the cart-pole simulation environment.